A listener initially 150 m from the source hears the sound with intensity I1. When the listener moved closer to the source half of its initial distance, how many times the sound intensity he hear increases?




1
Expert's answer
2021-02-17T10:56:24-0500

The intensity of the sound is inversely proportional to the distance from it. Thus:


"I_1 \propto\dfrac{1}{r_1^2}\\
I_2 \propto\dfrac{1}{r_2^2}"

where "I_1, I_2" are the initial and the final intensity respectively, "r_1,r_2" are the initial and the final distances. Thus, if "r_1 = 2r_2", the ration of intensities is:


"\dfrac{I_2}{I_1} = \dfrac{r_1^2}{r_2^2} = \dfrac{(2r_2)^2}{r_2^2} = 4"

The sound intensity increases 4 times.


Answer. 4.